The age of Ama is three more than kofi.If the product of their ages is 36years,find kofi's age 15.Ama's age 21.
Answer by Alan3354(69443)

The age of Ama is three more than kofi.If the product of their ages is 36years,find kofi's age 15.Ama's age 21.
--------------
15*21 = 315
----------------
A = K + 3
K*(K+3) = 36
K^2 + 3K - 36 = 0
Kofi is ~ 4.6847 years
Ama is ~ 7.6847 years